OpenStudy (anonymous):

Can anyone help me with Algebra 2 Simpifying?

OpenStudy (anonymous):

4(-2)(-(7-3)(5-2)2)

OpenStudy (austinl):

\(4(-2)(-(7-3)(5-2)2)\) What is 7-3? 5-2?

OpenStudy (anonymous):

4 and 5 right?

OpenStudy (austinl):

5-2=3 So now we have, \(4(-2)~(-(5)(3)2)\) What is \(5\times3\times2\)? What is \(4\times -2\)?

OpenStudy (anonymous):

30 and 2?

OpenStudy (austinl):

\(4\times-2=-8\) And we also have -30, so, \(-8\times-30=?\)

OpenStudy (anonymous):

192?

OpenStudy (austinl):

\(8\times30\ne192\) Do you have a calculator?

OpenStudy (anonymous):

no

OpenStudy (austinl):

Ok, here is a trick, What is \(8\times3\)?

OpenStudy (anonymous):

24

OpenStudy (austinl):

Now put a zero on the end of that :)

OpenStudy (anonymous):

240?

OpenStudy (austinl):

Correct!

OpenStudy (austinl):

\(8\times 30=240\)

OpenStudy (anonymous):

thank you so much:)

OpenStudy (austinl):

You are very welcome!